# Language Models as Critical Thinking Tools: A Case Study of Philosophers

### v1

@@ 7 Conclusion: Towards Living Script
In his masterwork Jerusalem, Moses Mendelssohn writes that philosophy has too long prioritized a dead form of interaction, one which stifles human interaction and innovation: “We teach and instruct one another only through writings; we learn to know nature and human only from writings. We work and relax, edify and amuse ourselves through scribbling…” (Mendelssohn, 1983, 41). In response, Mendelssohn calls for a turn towards a living script, “arousing the mind and heart, full of meaning, continuously inspiring thought”. The living script is a way of engaging with tools that inspire and support our critical thinking; it is an ideal both for LM researchers, philosophers, and all of us – as thinkers and humans – to aspire towards. As technologies for reading and writing our living script, LMs can offer critical thinkers a more wide and accessible set of ways to support individuals in their development of ideas and to shape disciplinary practices and cultures. As for the rights and responsibilities to think – we should respond, with John: “We claim them all.”
@@
Although exploring ‘uncommon sense’ is important for critical thinking, we acknowledge that it can also be a deeply uncomfortable and unsettling experience. Disagreement can feel awkward in many contexts in daily life, even though it may not in designated spaces: “one of the best gifts a philosopher can give another is a good counterexample… in philosophy, we like a challenge, a pushback, for people to think that we’re wrong. That’s where philosophers thrive” (P5). Moreover, common sense encodes certain ethical or moral norms, such as pain is bad and racism is unjust; critical thinking tools may facilitate the revisiting and challenging of these norms in apparently inappropriate ways. To be sure, there is great value in this practice. We may not only to believe in true things but also know the right or best reasons for why we should believe in them (why is racism bad?), since having poor reasons for a belief may undermine the belief. Moreover, supposedly obvious moral principles and norms can be utilized to support unsavory positions (e.g., racism is bad, so we should only pursue ‘colorblind’ public policy); it is difficult to identify this if one does not adopt a critical view towards the entire system. Nevertheless, LMs can serve many purposes, and being critical thinking tools is just one of them; low-selfhood and low-initiative tools are needed to accomplish many other tasks. Users should consent to critical interactions with LMs.
@@
Some interviewees expressed that LMs raised difficult questions about academic integrity and authorship of ideas. It should be noted that because critical thinking tools are intended to support the process of thinking rather than replacing it, there is little risk of outright plagiarism, provided the tools are designed properly and used as intended. Nevertheless, there are interesting ethical questions about ownership of ideas with respect to involvement in their development. If a colleague’s offhand comment sparks an idea, leading to a publication, (how) should the colleague be credited? What if instead they intentionally discuss and develop an idea with you? This is an ongoing conversation over how LMs as critical thinking tools should be credited, but there seems to be little basis for stopping the development of LMs as critical thinking tools because of ethical concerns of authorship.
Acknowledgement
We greatly appreciate each of the philosophers who took time and care to talk with us.
Appendix A Appendix
A.1 Interview Questions and Guidelines

### v2

@@ 7 Conclusion: Towards Living Script
In his masterwork Jerusalem, Moses Mendelssohn writes that philosophy has too long prioritized a dead form of interaction, one which stifles human interaction and innovation: “We teach and instruct one another only through writings; we learn to know nature and human only from writings. We work and relax, edify and amuse ourselves through scribbling…” (Mendelssohn, 1783, 41). In response, Mendelssohn calls for a turn towards a living script, “arousing the mind and heart, full of meaning, continuously inspiring thought”. The living script is a way of engaging with tools that inspire and support our critical thinking; it is an ideal both for LM researchers, philosophers, and all of us — as thinkers and humans — to aspire towards. As potential technologies for reading and writing our living script, LMs can offer critical thinkers a more wide and accessible set of ways to support the development of ideas and to shape disciplinary practices and cultures. In the face of intellectual automation, it begins by saying, with John, for the rights and responsibilities to critically think: “We claim them all.”
@@
Although exploring ‘uncommon sense’ is important for critical thinking, we acknowledge that it can also be a deeply uncomfortable and unsettling experience. Disagreement can feel awkward in many contexts in daily life, even though it may not in designated spaces: “one of the best gifts a philosopher can give another is a good counterexample… in philosophy, we like a challenge, a pushback, for people to think that we’re wrong. That’s where philosophers thrive” (P5). Moreover, common sense encodes certain ethical or moral norms, such as “pain is bad” and “racism is unjust”; critical thinking tools may facilitate the revisiting and challenging of these norms in apparently inappropriate ways. To be sure, there is great value in this practice. We may not only want to believe in true things but also know the right or best reasons for why we should believe in them (in what sense of ‘bad’ is pain bad? why is racism unjust?), since having poor reasons for a belief may undermine the belief without our knowledge. Moreover, supposedly obvious moral principles and norms can be utilized to support positions we might think to be unsavory or misguided (e.g., racism is unjust, so we should only pursue a strictly ‘colorblind’ public policy); it is difficult to identify this if one does not adopt a critical view towards the entire system. Nevertheless, LMs can serve many purposes, and being critical thinking tools is just one of them. Low-selfhood and low-initiative tools are needed to accomplish many other important tasks. Users should consent to critical interactions with LMs.
@@
Some interviewees expressed that LMs raised difficult questions about academic integrity and authorship of ideas. It should be noted that because critical thinking tools are intended to support the process of thinking rather than replacing it, there is little risk of outright plagiarism, provided the tools are designed properly and used as intended. Nevertheless, there are interesting ethical questions about ownership of ideas with respect to involvement in their development. If a colleague’s offhand comment sparks an idea, leading to a publication, (how) should the colleague be credited? What if instead they intentionally discuss and develop an idea with you? What is an author (Foucault, 1969b)? The question of how LMs as critical thinking tools should be credited joins the broader existing rich discourse of how generative AI in general should be credited in intellectual production (Hullman et al., 2023; Jenkins & Lin, 2023; Simon et al., 2024; Springer, 2024, inter alia).
Appendix A Interviewee Information Sheet
Table 1 provides high-level information about each interviewee which may be relevant to interpreting and contextualizing their views. The General Interest(s) feature describes the broad fields that the interviewees work in. The Notable Specific Interest(s) feature describes any specific topics in the field(s) mentioned in the General Interest(s) feature that the interviewees focus their work on. This feature is not exclusive, meaning that interviewees may also work on other topics outside of the specific interests. If the value for this feature is blank, then the interviewee’s work is sufficiently characterized by the value in the General Interest(s) feature. The Experience with LMs feature describes three levels of experience with using LMs: little to none, limited, and extensive. If interviewees have either limited or extensive experience with using LMs, the Uses of LMs feature describes their primary use: for teaching (e.g., using LMs to teach material, trying to understand features of LM-generated student submissions), for personal use (e.g., to improve productivity, for entertainment), for exploration (i.e., playing around with the LM out of curiosity to understand the technology better), and for research (i.e., their research is on LMs). Note that the following interviewees have published at least one article on some aspect of AI: (P5, P6, P13, P14).
ID Title General Interest(s) Notable Specific Interest(s) Experience with LMs Use of LMs
P1 Associate Professor Ethics, Political Philosophy Bioethics, Feminist Ethics Limited For teaching
P2 Associate Professor Philosophy of Science Philosophy of Biology Limited For exploration
P3 Professor Ethics, Aesthetics Meta-ethics Limited For teaching
P4 Professor Ethics, Political Philosophy Limited For personal use
P5 Assistant Professor Ethics Virtue ethics Limited For teaching
P6 Assistant Professor Ethics, Political Philosophy Philosophy of Technology, AI Extensive For research
P7 Assistant Professor Philosophy of Science Philosophy of Physics Extensive For personal use
P8 Associate Professor History of Philosophy German philosophy Limited For personal use
P9 Professor Philosophy of Science Extensive For exploration
P10 Associate Professor Ethics, History of Philosophy Philosophy of Technology Little to None
P11 Professor Philosophy of Science Philosophy of Statistics Little to None
P12 Professor Philosophy of Science Psychology Limited For exploration
P13 Associate Professor Philosophy of Science Philosophy of Biology Limited For class
P14 Professor Logic, Philosophy of Mind Semantics, Linguistics Extensive For exploration
P15 Assistant Professor Aesthetics Value theory, Literature Limited For exploration
P16 Professor Ethics, Political Philosophy Public and Global Policy Limited For exploration
P17 Teaching Professor Pedagogy, Epistemology Extensive For exploration
P18 Associate Professor Philosophy of Science Philosophy of Physics Limited For class
P19 Assistant Professor Ethics Moral psychology Limited For personal use
P20 Professor History of Philosophy Little to None
P21 Associate Professor Ethics Bioethics Little to None
Table 1: Interviewee information.
Appendix B Interview Questions and Guidelines